find the value of r so that the line that passes through each pair of the points has the given slope.
1) (r,2) (5,r), m=0

Answer by stanbon(75887)

find the value of r so that the line that passes through each pair of the points has the given slope.
1) (r,2) (5,r), m=0
-------------
slope = (r-2)/(5-r) = 0
r-2 = 0
r = 2
============================
Cheers,
Stan H.